answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What is the probability of rolling a 6-sided dice and getting an odd number three times and a row?

The cube has 6 faces, so it can come up in 6 different ways.

Three of the possible outcomes are odd numbers. (1, 3, and 5)

Probability of an odd number on the first roll = 3/6 = 1/2 .

Probability of an odd number on the second roll = 3/6 = 1/2 .

Probability of an odd number on the third roll = 3/6 = 1/2 .

Probability of an odd number on all 3 rolls = (1/2) x (1/2) x (1/2) = 1/8 = 0.125 = 12.5%

What numbers in the lottery have the most probability of winning?

They all have the same probability, even though in the past some have been drawn more often than others. It's still completely random though!

The probability of drawing somethingout of a bag?

If you reach into the bag, grasp, and withdraw your hand,

the probability of pulling something out is about 100%.

What are the odds of getting mauled by a bear?

You stand a better chance of being struck by lighting than mauled by a bear. According the Ontario Ministry of Natural Resources there are about 10 to 20 bear attacks in North America every year, with about 2 or 3 deaths.

Not sure what the odds are, but given the huge number of people that go tromping through the forest and the close proximity man and bear have to one another, the odds are pretty low.

List all the possible outcomes when the die is tossed once?

Th dice have 6 sides and each side have one number from 1-6 so that there are 6 possible outcomes if you rolled the dice

If a box contains one red ball and one blue ball and two people touch one ball randomly with a one in two chance of picking either what is the chance of the blue ball being picked twice?

One in four.

The possibilities are;

Red Red

Blue Red

Red Blue

Blue Blue

That means the number of ways an event can occur is one, while the number of possible outcomes is four. The definition of probability is the number of ways the event can occur over the number of possible outcomes, hence 1/4

The number of dots on a regular die?

There are six faces numbered 1 to 6 so:

6 + 5 + 4 + 3 + 2 + 1 = 21

How do you calculate a continuous joint probability distribution?

You simply have a function with two (or more) arguments which are continuous.

For example, z = p(x,y) would be a surface in 3 dimensions where x and y are the values taken by the two variables, X and Y respectively, and z is the probability associated with them.

w = g(x,y,z) would be a hyper-surface in 4-dimensional space and so on.

What amount of sampling error is acceptable in a national sample of 1500 to 3000 respondents?

There can be no set value.

An acceptable level of sampling error for a company making high precision machine parts is likely to be very different from the sampling error for household incomes, for example.

What is the probability of tossing a 2 or 3 on a number cube?

Each face has probability 1/6, since it's one face out of the possible 6.

So the probability of either of two faces is 2/6, or 1/3.

You roll 2 dice What is the probability that the sum of the dice is greater than 6 and 1 die shows a 3?

Each cube can land in 6 different ways.

Two cubes can land in (6 x 6) = 36 different ways.

There are 6 ways that succeed:

3 - 4

4 - 3

3 - 5

5 - 3

3 - 6

6 - 3

The probability is (6/36) = 1/6 = 16 and 2/3 percent

If the paired difference's are normal in a test of mean difference's then the distribution used for testing is the?

If the paired differences are normal in a test of mean differences, then the distribution used for testing is the

What is a sample space diagram in probabiliy?

It is a diagram of all possible outcomes of a probability experiment.

Where is the center of gravity in a structure?

For a structure (usually a tower) that is perpendicular to the ground, if the center of gravity were to go a distance over its base, it will fall over. Can be seen in the leaning tower of piza. The reason it has not fallen over is because its center of gravity has not exceeded the length of its base.

What is the probability of tossing seven coins and getting 6 heads?

The probability of getting only one tails is (1/2)7. With seven permutations of which flip is the tails, this gives a probability of:

P(six heads in seven flips) = 7*(1/2)7 = 7/128

What is the probability of a coin landing on heads and a dice landing on an Evan number?

Coin landing of heads = 1/2 (either heads or tails)

Dice landing on even number = 1/2 (no matter how many faces there are on the dice unless the number of faces is odd, 6 sided=3 even sides/6)

If you roll a nine-sided die what is the probability you will roll a value greater than zero?

If zero is the number on one of the 9 sides, then the probability is 8/9 = 88.9% (rounded)

If zero is not the number on any one of the 9 sides, then the probability is 100%.